OpenStudy (jotopia34):

Calc II question please... Can someone please explain if the sequence cosn/n^4 converges or diverges? I know the formal definition of Squeeze but I came up with ---1

OpenStudy (anonymous):

cosine is stuck between -1 and 1, so it converges for sure

OpenStudy (jotopia34):

\[a_n = \frac{ cosn }{ n^4 }\] \[-1\le(cosn/n^4)\le1\]

OpenStudy (jotopia34):

but I have to show the squeeze theorem for the convergence but I am not getting a limit that =L

OpenStudy (anonymous):

try \[-\frac{1}{n^4}\leq \frac{\cos(n)}{n^4}\leq \frac{1}{n^4}\]

OpenStudy (jotopia34):

why is it over n^4

OpenStudy (anonymous):

since the left and right both go to zero, so does the middle one

OpenStudy (jotopia34):

Are the left and right parts (x1/n^4)? Is that why

OpenStudy (anonymous):

you wrote \(\frac{\cos(n)}{n^4}\) right?

OpenStudy (jotopia34):

yes

OpenStudy (anonymous):

that is why i wrote \[-\frac{1}{n^4}\leq \frac{\cos(n)}{n^4}\leq \frac{1}{n^4}\]

OpenStudy (anonymous):

\(-1\leq \cos(n)\leq 1\) divide both sides by \(n^4\)

OpenStudy (jotopia34):

why wouldn't I multiply each side, like to get rid of it in the middle

OpenStudy (anonymous):

you are trying to show \(\frac{\cos(n)}{n^4}\) converges

OpenStudy (anonymous):

the numerator is stuck between -1 and 1, so the whole thing is stuck between \(-\frac{1}{n^4}\) and \(\frac{1}{n^4}\)

OpenStudy (jotopia34):

yes, I am but I have to show what you did in terms of the definition of squeeze

OpenStudy (anonymous):

it is the squeeze, it is squeezed between to things that go to zero

OpenStudy (jotopia34):

okay so can something converge to zero?

OpenStudy (jotopia34):

do sin and cos always converge to 0? I thought they are alternating sequences

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{-1}{n^4}=0\] and \[\lim_{n\to \infty}\frac{1}{n^4}=0\] and \[\frac{\cos(n)}{n^4}\] is squeezed between them

OpenStudy (jotopia34):

I see that now for sure.

OpenStudy (anonymous):

it is true that \(\cos(n)\) has no limit as \(n\to \infty\) but you have \[\frac{\cos(n)}{n^4}\]

OpenStudy (jotopia34):

so what would be convergence of \[\sum_{n=1}^{\infty} \cos n/n^4\]

OpenStudy (anonymous):

again it converges

OpenStudy (jotopia34):

but wouldn't it be comparable to a p series with p=4>1

OpenStudy (anonymous):

you can compare it to \[\sum\frac{1}{n^4}\]which converges like mad

OpenStudy (jotopia34):

uh oh, p>1 converges?

OpenStudy (anonymous):

in fact it is absolutely convergent

OpenStudy (anonymous):

yes if \(p>1\) then \(\sum\frac{1}{n^p}\) converges

OpenStudy (jotopia34):

actually, I am getting this because I know abs convergence is true if absolute value of a_n converges, correct?

OpenStudy (jotopia34):

so if theres no negative one to cos(n), then it would have to converge....is that right?

OpenStudy (anonymous):

yes

OpenStudy (jotopia34):

yayyyy!! thank you. you are awesome!!

OpenStudy (anonymous):

although in this case i think you have to use the comparison test ratio or root won't work

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

in fact not only does it converge, but it is rather small i think you can check here http://www.wolframalpha.com/input/?i=sum+1+to+infinity+cos%28n%29%2Fn^4

OpenStudy (jotopia34):

in the case of cos n x 1/n^4, is a_n less than b_n for all n?

OpenStudy (jotopia34):

and how can you tell?

OpenStudy (anonymous):

\[|\frac{\cos(n)}{n^4}|\leq \frac{1}{n^4}\]

OpenStudy (anonymous):

gotta run, good luck

OpenStudy (jotopia34):

thanks tons of help